# CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection

### v1

@@ Abstract
Sparse Autoencoders (SAEs) can extract interpretable features from large language models (LLMs) without supervision. However, their effectiveness in downstream steering tasks is limited by the requirement for contrastive datasets or large activation storage. To address these limitations, we propose CorrSteer, which selects features by correlating sample correctness with SAE activations from generated tokens at inference time. This approach uses only inference-time activations to extract more relevant features, thereby avoiding spurious correlations. It also obtains steering coefficients from average activations, automating the entire pipeline. Our method shows improved task performance on QA, bias mitigation, jailbreaking prevention, and reasoning benchmarks on Gemma 2 2B and LLaMA 3.1 8B, notably achieving a +4.1% improvement in MMLU performance and a +22.9% improvement in HarmBench with only 4000 samples. Selected features demonstrate semantically meaningful patterns aligned with each task’s requirements, revealing the underlying capabilities that drive performance. Our work establishes correlation-based selection as an effective and scalable approach for automated SAE steering across language model applications.
CorrSteer: Steering Improves Task Performance and Safety in LLMs through Correlation-based Sparse Autoencoder Feature Selection
Seonglae Cho1,2  Zekun Wu1,2  Adriano Koshiyama1,2 1Holistic AI  2University College London
@@ 1 Introduction
Sparse Autoencoders (SAEs) have emerged as a powerful tool for decomposing superposed representations in large language models (LLMs) into interpretable sparse latent dimensions (Huben et al., 2023). By reconstructing neural activations through a sparse bottleneck, SAEs effectively disentangle semantic features that can be leveraged for downstream tasks such as probing and steering (Bricken et al., 2023).
However, existing SAE-based steering approaches face significant limitations: (1) contrastive datasets (Soo et al., 2025) or large activation storage (Zhao et al., 2025; Arad et al., 2025) are required to identify the direction of the steering, and (2) they rely on the hidden states of context tokens to select both the features and their coefficients.
Figure 1: Top correlated features with MMLU in each layer of Gemma 2 2B.
Consequently, current use cases of SAE-based steering have been restricted to specific applications, such as bias mitigation (Durmus et al., 2024), knowledge unlearning (Muhamed et al., 2025; Wang et al., 2025; Zhou et al., 2025; Cywiński and Deja, 2025), and jailbreaking prevention (O’Brien et al., 2025). Moreover, SAE feature selection in these applications does not directly reflect language models’ generation capabilities, potentially limiting their applicability.
Figure 2: Benchmark performance of CorrSteer variants compared with the baseline on Gemma 2 2B.
To address these limitations, this work introduces CorrSteer, which leverages generation-time features by correlating with task outcomes for task-specific feature selection and steering coefficient determination. Our approach employs Pearson correlation, which captures linear relationships, a lightweight yet effective criterion for rapidly identifying task-relevant features from minimal samples. Focusing on steering static behaviors, CorrSteer’s effectiveness is demonstrated on generation tasks by improving benchmark accuracy on MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024), BBQ (Parrish et al., 2022), HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024), and SimpleQA (Wei et al., 2024). Finally, by defining SER (Side Effect Ratio), three variants of CorrSteer are compared targeting the minimization of SER against fine-tuning.
2 Background
@@
2.1 Sparse Autoencoders
Sparse Autoencoders (Huben et al., 2023; Bricken et al., 2023) address the superposition problem by learning to decompose neural activations into interpretable, sparse features. Given an activation vector $\mathbf{x}\in\mathbb{R}^{d}$, an SAE learns an encoder $f_{\text{enc}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and decoder $f_{\text{dec}}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{d}$ where $k\gg d$, such that:
$\mathbf{z}$ $=f_{\text{enc}}(\mathbf{x})=\text{Activation}(\mathbf{W}_{\text{enc}}\mathbf{x}+\mathbf{b}_{\text{enc}})$ (1)
$\hat{\mathbf{x}}$ $=f_{\text{dec}}(\mathbf{z})=\mathbf{W}_{\text{dec}}\mathbf{z}+\mathbf{b}_{\text{dec}}$ (2)
The training objective is usually a combination of reconstruction loss with sparsity regularization:
$\mathcal{L}=\|\mathbf{x}-\hat{\mathbf{x}}\|^{2}+\lambda\|\mathbf{z}\|_{1}$ (3)
2.2 Steering Vectors
@@
Steering vectors (Subramani et al., 2022) represent a class of methods for controlling neural network outputs by manipulating internal activations. Traditional approaches, such as CAA (Rimsky et al., 2024), compute activation differences between contrasting examples and apply these differences. For precise control to not inadvertently affect related behaviors simultaneously, PaCE (Luo et al., 2024) utilizes sparse coding for orthogonal steering directions.
2.3 SAE-based Steering
@@
SAE-based steering leverages SAE latents for predictable control based on feature semantics. SAE-TS (Chalnev et al., 2024; Soo et al., 2025) reduces the side effects of steering by linearly approximating feature directions. SPARE (Zhao et al., 2025) utilizes Mutual Information to select features and their coefficients but requires large activation storage due to its non-linearity. DSG (Muhamed et al., 2025) utilizes Fisher Information Matrix to select features but requires contrastive datasets and additional backward computation. Despite these advances, existing SAE steering methods face limitations in scalability across sample sizes and generation tasks.
3 Method
Linear correlation offers both interpretability and faithfulness as a criterion for feature selection. SAEs capture linear relationships, consistent with the Linear Representation Hypothesis (Socher et al., 2013; Faruqui et al., 2015; Park et al., 2023), and have a proven ability to disentangle interpretable features in a linear manner. The faithfulness of Pearson correlation is further supported by recent work from Oikarinen et al. (2025).
3.1 Correlation-based Feature Selection
Our approach, CorrSteer, centers on the observation that features most correlated with task performance are likely to be relevant for steering. The approach employs the Pearson correlation coefficient, applied only to generation-time features—specifically to the last token at each step.
@@ 3.1 Correlation-based Feature Selection
Given a set of SAE features $\mathbf{z}=[z_{1},z_{2},\ldots,z_{D}]$ and corresponding task performance scores $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$ for $n$ samples, the correlation for each feature $i$ is computed as:
@@
To handle the computational challenges of large SAE feature dictionaries (typically $10^{4}$$10^{5}$ features), a streaming correlation accumulator is implemented that maintains $O(1)$ memory complexity:
@@
  $n\leftarrow n+\text{batch\_size}$
@@
 Compute correlations:
$medmath{r_{i}=\frac{n\sum x_{i}y_{i}-\sum x_{i}\sum y_{i}}{\sqrt{(n\sum x_{i}^{2}-(\sum x_{i})^{2})(n\sum y_{i}^{2}-(\sum y_{i})^{2})}}}$
@@
This computation is $O(1)$ with respect to sample size, and $O(LD)$ for fixed layer count $L$ and SAE latent dimension $D$. For generation tasks requiring multiple tokens, max-pooling is employed over valid token positions to aggregate feature activations, as empirically validated in our pooling comparison study (Table 5).
3.2 Coefficient Calculation
The steering coefficient for a selected feature $i$ is computed as the mean activation value among samples where task performance is positive. This approach is preferable to contrastive-based calculation, since SAE features produce non-negative activations due to their ReLU-based activation functions (Bricken et al., 2023) and thus cannot be meaningfully subtracted in a contrastive manner; negative activations are often unrelated noise (Joseph Bloom, 2024).
$c_{i}=\frac{1}{|\{j:y_{j}>0\}|}\sum_{j:y_{j}>0}z_{i,j}$ (5)
This ensures that the steering magnitude reflects the natural activation scale of the feature during successful task performance.
3.3 Steering Implementation
During inference, steering is applied by modifying residual stream activations. For a selected feature $i$, with coefficient $c_{i}$ and SAE decoder weights $\mathbf{W}_{\text{dec}}$ (the feature direction (Templeton et al., 2024)), the steering vector is:
$\mathbf{v}_{\text{steer}}=c_{i}\cdot\mathbf{W}_{\text{dec}}[:,i]$ (6)
The modified activation is:
$\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{v}_{\text{steer}}$ (7)
Steering is applied to tokens corresponding to the positions from which the features were originally extracted, rather than only to the last token (Luo et al., 2024; Rimsky et al., 2024) or every token (Soo et al., 2025).
3.4 Feature Extraction Strategies
Figure 3: Comparison of selected features between CorrSteer-1, CorrSteer-A, and CorrSteer-P on the BBQ disambiguous task across all layers of Gemma 2 2B. Red points indicate features selected by each method.
@@ 3.4 Feature Extraction Strategies
For each layer $\ell$, we obtain SAE activations from the residual stream and rank features by correlation with task performance. The method compares a global view aggregated across layers and a layer-wise view for selecting features to steer. Three strategies are implemented:
  • CorrSteer-1: Select the single highest-correlated feature from the global view, allowing cross-layer feature competition.
  • CorrSteer-A: Select the top-1 feature within each layer.
  • CorrSteer-P: Prune features from CorrSteer-A using a validation set, retaining only those that improve baseline performance to reduce side effects from spurious correlations.
Each method has distinct pros and cons, discussed in Section 5. Only positively correlated features are selected to ensure steering induces positive activation, as our ablation study confirms that negative correlation features consistently degrade performance 6. All methods are fully automated based on observed activations without hyperparameter tuning.
3.5 Side Effect Ratio
A key challenge in correlation-based feature selection is distinguishing features that causally contribute to task success from those that merely correlate due to the model’s internal state, potentially causing unintended side effects. To quantify side effects, the Side Effect Ratio (SER) is defined as the proportion of negatively changed answers among all changed answers:
$SER=\frac{\text{\# negatively changed answers}}{\text{\# all changed answers}}$ (8)
@@ 3.5 Side Effect Ratio
This measure does not isolate the side effect of each individual feature; rather, it serves as a combined metric reflecting how well the selected features are optimized for the task without degrading the model’s original abilities. To reduce side effects, the approach focuses on features activated during the generation process, under the hypothesis that generation-time activations are more likely to be causally relevant to output. This inference-time focus is empirically validated by our pooling experiments (Table 5). Additionally, in the multi-layer approach, a validation-based filtering mechanism is introduced (CorrSteer-P), retaining only features that demonstrate actual steering effectiveness.
4 Experiments
Figure 4: Relation between consumed sample counts and test performance, final matched count of the selected features, and the most correlated features from each layer of Gemma 2 2B. The dotted lines show the baseline of default LLM performance and constrained decoding performance on MMLU answer options.
4.1 Experimental Setup
CorrSteer is evaluated across diverse generation benchmarks to demonstrate practical effectiveness.
@@ 4.1 Experimental Setup
Models and SAEs: Experiments are conducted using Gemma-2 2B (Team, 2024a) and LLaMA-3.1 8B (Team, 2024b) models, paired with their corresponding SAE releases from Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), respectively. Both SAE families employ JumpReLU activation (Rajamanoharan et al., 2024). Additionally, the Gemma-2-2B-IT model with SAEs is employed, leveraging the fact that SAEs are typically transferable across fine-tuned models (Kissane et al., 2024), with proven low loss reported in the Gemma Scope paper (Lieberum et al., 2024).
Datasets: Our evaluation encompasses multiple benchmark categories:
  • Question Answering: MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024)
  • Reasoning: GSM8k (Cobbe et al., 2021)
  • Bias Evaluation: BBQ (Parrish et al., 2022)
  • Factuality: SimpleQA (Wei et al., 2024)
@@
  • Safety: HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024)

### v2

# CorrSteer: Generation-Time LLM Steering via Correlated Sparse Autoencoder Features

@@ Abstract
Sparse Autoencoders (SAEs) can extract interpretable features from large language models (LLMs) without supervision. However, their effectiveness in downstream steering tasks is limited by the requirement for contrastive datasets or large activation storage. To address these limitations, we propose CorrSteer, which selects features by correlating sample correctness with SAE activations from generated tokens at inference time. This approach uses only inference-time activations to extract more relevant features, thereby reducing spurious correlations. It also obtains steering coefficients from average activations, automating the entire pipeline. Our method shows improved task performance on QA, bias mitigation, jailbreaking prevention, and reasoning benchmarks on Gemma-2 2B and LLaMA-3.1 8B, notably achieving a +3.3% improvement in MMLU performance with 4000 samples and a +27.2% improvement in HarmBench with only 108 samples. Selected features demonstrate semantically meaningful patterns aligned with each task’s requirements, revealing the underlying capabilities that drive performance. Our work establishes correlation-based selection as an effective and scalable approach for automated SAE steering across language model applications.
Figure 1: System diagram of CorrSteer. CorrSteer selects task-relevant SAE features by correlating generated-token activations with outcomes, and constructs steering vectors applied as CorrSteer-S, CorrSteer-A, or CorrSteer-P. This generation-time steering shifts model behavior from unintended to intended responses while reducing side effects.
@@ 1 Introduction
Sparse Autoencoders (SAEs) have emerged as a powerful tool for decomposing superposed representations in large language models (LLMs) into interpretable sparse latent dimensions (Huben et al., 2023). By reconstructing neural activations through a sparse bottleneck, SAEs disentangle semantic features that can be leveraged for downstream tasks such as probing and steering (Bricken et al., 2023). However, existing SAE-based steering approaches face significant limitations: (1) contrastive datasets (Soo et al., 2025) or large activation storage (Zhao et al., 2025; Arad et al., 2025) are required to identify the direction of the steering, and (2) they rely on the hidden states of context tokens to select both the features and their coefficients. Consequently, current use cases of SAE-based steering have been restricted to specific applications, such as bias mitigation (Durmus et al., 2024), knowledge unlearning (Muhamed et al., 2025; Wang et al., 2025; Zhou et al., 2025; Cywiński & Deja, 2025), and jailbreaking prevention (O’Brien et al., 2025). Moreover, SAE feature selection in these applications does not directly reflect language models’ generation capabilities, potentially limiting their applicability. To address these limitations, this work introduces CorrSteer, which leverages generation-time features by correlating with task outcomes for task-specific feature selection and steering coefficient determination. Our approach employs Pearson correlation, which captures linear relationships, a lightweight yet effective criterion for rapidly identifying task-relevant features from minimal samples. Focusing on steering static behaviors, CorrSteer’s effectiveness is demonstrated on generation tasks by improving benchmark accuracy on MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024), BBQ (Parrish et al., 2022), HarmBench (Mazeika et al., 2024), XSTest (Röttger et al., 2024), and SimpleQA (Wei et al., 2024). Finally, by defining SER (Side Effect Ratio), three variants of CorrSteer are compared targeting the minimization of SER against fine-tuning.
2 Related Work
@@
Steering Vectors (Subramani et al., 2022) represent a class of methods for controlling neural network outputs by manipulating internal activations. Traditional approaches, such as CAA (Rimsky et al., 2024), compute activation differences between contrasting examples and apply these differences. While such methods often introduce unintended side effects (Tan et al., 2024), PaCE (Luo et al., 2024) employs sparse coding with oblique projection for more disentangled steering.
@@
SAE-based Steering leverages Sparse Autoencoder latents for predictable control based on feature semantics. SAE-TS (Chalnev et al., 2024; Soo et al., 2025) reduces the side effects of steering by linearly approximating feature directions. SPARE (Zhao et al., 2025) utilizes Mutual Information to select features and their coefficients but requires large activation storage due to its non-linearity. DSG (Muhamed et al., 2025) utilizes Fisher Information Matrix to select features but requires contrastive datasets and additional backward computation. Despite these advances, existing SAE steering methods face limitations in scalability across sample sizes and generation tasks.
Recent work has shown that SAEs capture linear relationships consistent with the Linear Representation Hypothesis (Socher et al., 2013; Faruqui et al., 2015; Park et al., 2023), and Pearson correlation has been demonstrated as a faithful measure for such linear dependencies (Oikarinen et al., 2025). These findings motivate our proposed approach, CorrSteer, which leverages correlation-based feature selection for automated and scalable SAE steering. This simplicity, combined with scalability and interpretability, distinguishes CorrSteer from prior SAE steering methods
3 The CorrSteer Method
Figure 1 provides an overview of CorrSteer, illustrating how correlation-based feature selection and steering interventions are applied. CorrSteer is a simple yet scalable pipeline that steers language models by linking generation-time SAE activations with task outcomes. Our method first identifies task-relevant features via correlation, then assigns coefficients from their natural activation scales, and finally applies steering vectors during inference. This design emphasizes three advantages over prior SAE-based steering: simplicity, scalability, and interpretability.
3.1 Correlation-Guided Feature Selection
The central idea of CorrSteer is that features most correlated with task performance are also the most promising candidates for steering. Pearson correlation is well-suited for SAE’s inherently linear architecture where features are designed to be linearly combined (Bricken et al., 2023), aligning with the Linear Representation Hypothesis (Park et al., 2023) and leveraging correlation as a faithful measure for linear dependencies in neural representations (Oikarinen et al., 2025). To capture this relationship, we compute correlations only on generation-time activations, focusing on the last generated token at each step, since these activations are most directly tied to model output correctness.
@@ 3.1 Correlation-Guided Feature Selection
Formally, given a set of SAE features $\mathbf{z}=[z_{1},z_{2},\ldots,z_{D}]$ and corresponding task performance scores $\mathbf{y}=[y_{1},y_{2},\ldots,y_{n}]$ for $n$ samples, the correlation for each feature $i$ is computed as:
@@
To handle the computational challenges of large SAE feature dictionaries (typically $10^{4}$-$10^{5}$ features), a streaming correlation accumulator is implemented that maintains $O(1)$ memory complexity:
@@
  $n\leftarrow n+|\mathbf{y}_{\text{batch}}|$
@@
 Compute correlations for each feature $i$:
$r_{i}=\frac{n\sum x_{i}y_{i}-\sum x_{i}\sum y_{i}}{\sqrt{(n\sum x_{i}^{2}-(\sum x_{i})^{2})(n\sum y_{i}^{2}-(\sum y_{i})^{2})}}$
@@
This computation maintains $O(1)$ space complexity with respect to sample size, while time complexity is $O(N)$ for $N$ samples, and $O(LD)$ for fixed layer count $L$ and SAE latent dimension $D$. For generation tasks requiring multiple tokens, max-pooling is employed over valid token positions to aggregate feature activations, as empirically validated in our pooling comparison study (Table 5).
3.2 Coefficient Estimation from Positive Outcomes
For each selected feature $i$, we define its steering coefficient as the mean activation over samples with positive task outcomes. Formally:
$c_{i}=\frac{1}{|\{j:y_{j}>0\}|}\sum_{j:y_{j}>0}z_{i,j}.$ (2)
This formulation directly anchors the steering magnitude to the feature’s natural activation scale during successful performance. Unlike contrastive-based methods, it leverages the non-negativity of SAE activations (arising from ReLU) (Bricken et al., 2023), thereby avoiding ill-posed subtraction between activation states and ensuring stable, semantically faithful steering. These coefficients are then used at inference time to construct steering vectors that modify the model’s residual stream.
3.3 Inference-Time Steering Mechanism
At inference time, steering modifies residual stream activations at each token generation step. For a selected feature $i$ with coefficient $c_{i}$ and SAE decoder weights $\mathbf{W}_{\text{dec}}$ (its feature direction (Templeton et al., 2024)), the steering vector $\mathbf{v}_{\text{steer}}=c_{i}\cdot\mathbf{W}_{\text{dec}}[:,i]$ is added to the residual stream as $\mathbf{x}^{\prime}=\mathbf{x}+\mathbf{v}_{\text{steer}}$ for each generated token position. We apply steering exclusively at token positions where the corresponding features were extracted, rather than uniformly across all tokens (Soo et al., 2025) or restricted to the final token (Luo et al., 2024; Rimsky et al., 2024). Since many benchmarks involve multi-token generations, this raises the question of how to aggregate activations across tokens when computing correlations and coefficients, which we address next.
3.4 Pooling Strategy for Feature Aggregation.
Two pooling strategies are available for coefficient and correlation calculations: max-pooling and mean-pooling. For multi-token generation tasks, max-pooling consistently outperforms mean-pooling, as empirically demonstrated in Table 5, likely due to its better capture of peak feature activations relevant to task success. However, for coefficient calculation in longer generation tasks such as GSM8K reasoning, mean-pooling is preferred as max-pooling produces excessively large coefficient values. Applying these large coefficients to every generated token degrades performance, leading to the adoption of mean-pooling for reasoning tasks. Having established how features are aggregated across tokens, we next turn to how features are selected across layers.
3.5 Automated Multi-Layer Feature Selection
@@
For each layer $\ell$, we extract SAE activations from the residual stream and rank features by their correlation with task performance. We consider both a global view aggregating correlations across layers and a layer-wise view that preserves layer-specific structure. Based on these perspectives, we implement three fully automated strategies (no hyperparameter tuning required):
  • CorrSteer-S. Select the single most positively correlated feature across all layers (global view). This minimal variant tests whether a single feature suffices for causal performance improvements.
  • CorrSteer-A. Select the top positively correlated feature from each layer. This design probes whether layer-wise features collectively form circuits that enhance task performance.
  • CorrSteer-P. Begin with CorrSteer-A and apply validation-based pruning, retaining only those features that improve over the non-steered model. This enables finer-grained subcircuit analysis.
Only positively correlated features are retained, as ablation experiments confirm that negatively correlated features consistently degrade performance (Appendix 6). Figure 2 illustrates these strategies on the BBQ (disambiguous) task across all layers of Gemma-2 2B, highlighting how CorrSteer-S, CorrSteer-A, and CorrSteer-P differ in terms of selected feature distribution (red points). While CorrSteer-S focuses on a single dominant signal, CorrSteer-A distributes selections across layers, and CorrSteer-P prunes this set to retain only features that yield improvements. These differences highlight distinct trade-offs in global versus layer-wise selection. However, feature selection may also introduce unintended side effects, which we address next.
Figure 2: Comparison of features selected by CorrSteer-S, CorrSteer-A, and CorrSteer-P on BBQ (disambiguous) across all Gemma-2 2B layers. Red points denote selected features.
3.6 Quantifying Side Effects via SER
Correlation-based feature selection risks capturing spurious associations rather than causal drivers, leading to unintended degradations. We quantify this with the Side Effect Ratio (SER), defined as
$\text{SER}=\frac{\#\text{ negatively changed answers}}{\#\text{ all changed answers}}.$ (3)
@@ 3.6 Quantifying Side Effects via SER
Lower SER indicates more reliable steering, isolating features that improve performance without harmful side effects. This measure does not isolate the side effect of each individual feature; rather, it serves as a combined metric reflecting how well selected features are optimized for the task without degrading the model’s original abilities. To reduce side effects, the approach focuses on features activated during generation, under the hypothesis that generation-time activations are more likely causally relevant to output. This inference-time focus is empirically validated by our pooling experiments (Table 5). Additionally, in the multi-layer approach, a validation-based filtering mechanism is introduced (CorrSteer-P), retaining only features that demonstrate steering effectiveness.
4 Experimental Setup
@@
Experiments are conducted using Gemma-2 2B (Team, 2024a) and LLaMA-3.1 8B (Team, 2024b) models, paired with their corresponding SAE releases from Gemma Scope (Lieberum et al., 2024) and LLaMA Scope (He et al., 2024), respectively. Both SAE families employ JumpReLU activation (Rajamanoharan et al., 2024). Additionally, the Gemma-2 2B-IT model with SAEs is employed, leveraging the fact that SAEs are typically transferable across fine-tuned models (Kissane et al., 2024), with proven low loss reported in the Gemma Scope paper (Lieberum et al., 2024).
Evaluation Benchmarks We evaluate CorrSteer on a suite of benchmarks spanning five categories:
  • Knowledge: MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024) test broad-domain expertise under zero-shot settings.
  • Reasoning: GSM8K (Cobbe et al., 2021) probes multi-step mathematical reasoning ability.
  • Bias: BBQ (Parrish et al., 2022) measures sensitivity to social bias and stereotypes.
  • Factuality: SimpleQA (Wei et al., 2024) assesses short-form factual consistency.
@@
  • Safety: HarmBench (Mazeika et al., 2024) and XSTest (Röttger et al., 2024) evaluate resistance to unsafe or sensitive content generation.